OpenStudy (anonymous):

5.Griselda budgets forty dollars per month before sales tax for clothes. How much will she spend on clothes each month if she is charged seven percent sales tax? thirty-seven dollars and twenty cents forty-two dollars and eighty cents forty-seven dollars

OpenStudy (anonymous):

ok i should learn to read she budgets $40 before tax then increase by 7% by finding \[40\times (1.07)\]

OpenStudy (anonymous):

7% of $40

OpenStudy (anonymous):

\(.07\times 40=.7\times 4=2.8\) so \(42.80\)

OpenStudy (anonymous):

Thats what I just did. I got it right!!!! thx